OpenStudy (anonymous):

If f(0) = -2 and f(x + 2) - f(x) = 5, and f(x) is linear, what is the equation for f(x)?

OpenStudy (anonymous):

We know: \[ y=ax+b\\ -2=a(0)+b\\ b=-2 \] So:\[ y=a(x+2)-2-(ax-2)=ax+2a-2-ax+2=2a=5\\ a=\frac{5}{2} \] So, we are done.